# Mixture of Knowledge Minigraph Agents for Literature Review Generation

### v2

@@ Abstract
Literature reviews play a crucial role in scientific research for understanding the current state of research, identifying gaps, and guiding future studies on specific topics. However, the process of conducting a comprehensive literature review is yet time-consuming. This paper proposes a novel framework, collaborative knowledge minigraph agents (CKMAs)111https://aaai-2025-4471.github.io/, to automate scholarly literature reviews. A novel prompt-based algorithm, the knowledge minigraph construction agent (KMCA), is designed to identify relations between concepts from academic literature and automatically constructs knowledge minigraphs. By leveraging the capabilities of large language models on constructed knowledge minigraphs, the multiple path summarization agent (MPSA) efficiently organizes concepts and relations from different viewpoints to generate literature review paragraphs. We evaluate CKMAs on three benchmark datasets. Experimental results show the effectiveness of the proposed method, further revealing promising applications of LLMs in scientific research.
@@ Introduction
To handle these complex relations, we propose equipping LLMs with structural knowledge. Knowledge graphs, which represent entities as nodes and their relations as edges, are potential solutions. However, it is challenging to find a general-purpose knowledge graph that covers relations in literature reviews. Instead of having a single graph for everything (Narayanan et al. 2017), we propose knowledge minigraphs for reference documents of interest. Knowledge minigraphs are small-scale graphs that comprise concepts dynamically extracted from reference documents as nodes and their relations as edges. They omit detailed documentation and highlight relations between concepts.
@@
To generate a summary that is representative of the overall content, graph-based methods construct external graphs to assist document representation and cross-document relation modeling, achieving promising progress. In this regard, LexRank (Erkan and Radev 2004) and TextRank (Mihalcea and Tarau 2004) first introduced graphs to extractive text summarization in 2004. They compute sentence importance using a graph representation of sentences to extract salient textual units from documents as summarization. In 2020, Wang et al. propose to extract salient textual units from documents as summarization using a heterogeneous graph consisting of semantic nodes at several granularity levels of documents (Wang et al. 2020). In 2022, Wang et al. incorporate knowledge graphs into document encoding and decoding, generating the summary from a knowledge graph template to achieve state-of-the-art performance (Wang et al. 2022). However, to the best of our knowledge, no existing work integrates LLMs into graph-based methods to leverage their natural language understanding capabilities for improved graph construction and summary generation.
@@
In the first iteration, we apply the knowledge minigraph prompt to the first chunk of reference papers to derive the initial knowledge minigraph . For subsequent iterations, we transform the intermediate knowledge minigraph into a text representation . This text representation is then input along with the -th chunk of reference papers , allowing for dynamic updates to derive knowledge minigraph .
@@
| Proposed | 36.41 | 8.78 | 34.16 | 6.22 | 32.31 | 5.36 | |
Summarization router We design a router to evaluate different experts’ summaries and automatically select the most desirable summary as the final output. Without requiring additional side information, this paper proposes a self-evaluation strategy. In detail, we observe that there are agreements between the experts’ viewpoints and their generated summaries. We propose to quantify the degree of agreement for each summary using the ROUGE-1 score (Lin 2004), which measures the overlap between a generated summary and other summaries. We then select the summary with the highest degree of agreement, which indicates that its understanding has the highest likelihood of being supported by other experts, or in other words, is relatively more acceptable. Mathematically, the final summary can be denoted as:
@@
This section presents ablation studies to investigate the performance gains brought by the designed modules in CKMAs. We first ablate the knowledge minigraph construction agent (KMCA) and the multiple path summarization agent (MPSA), respectively. When ablating KMCA, we no longer construct the knowledge minigraph and do not include it as part of the MPSA instruction. When removing MPSA, we use the instruction in Table 1 to directly generate a single summary as the final summary. We report the performance of the ablated version of the model in the first line of Table 3.

### v3

@@ Abstract
Literature reviews play a crucial role in scientific research for understanding the current state of research, identifying gaps, and guiding future studies on specific topics. However, the process of conducting a comprehensive literature review is yet time-consuming. This paper proposes a novel framework, collaborative knowledge minigraph agents (CKMAs), to automate scholarly literature reviews. A novel prompt-based algorithm, the knowledge minigraph construction agent (KMCA), is designed to identify relations between concepts from academic literature and automatically constructs knowledge minigraphs. By leveraging the capabilities of large language models on constructed knowledge minigraphs, the multiple path summarization agent (MPSA) efficiently organizes concepts and relations from different viewpoints to generate literature review paragraphs. We evaluate CKMAs on three benchmark datasets. Experimental results show the effectiveness of the proposed method, further revealing promising applications of LLMs in scientific research.
Project — https://minigraph-agents.github.io/
@@ Introduction
To handle these complex relations, we propose equipping LLMs with structural knowledge. Knowledge graphs, which represent entities as nodes and their relations as edges, are potential solutions. However, it is challenging to find a general-purpose knowledge graph that covers relations in various literature reviews. Instead of having a single graph for everything (Narayanan et al. 2017), we propose knowledge minigraphs for reference documents of interest. Knowledge minigraphs are small-scale graphs that comprise concepts dynamically extracted from reference documents as nodes and their relations as edges. They omit detailed documentation and highlight relations between concepts.
@@
To generate a summary that is representative of the overall content, graph-based methods construct external graphs to assist document representation and cross-document relation modeling, achieving promising progress. In this regard, LexRank (Erkan and Radev 2004) and TextRank (Mihalcea and Tarau 2004) first introduce graphs to extractive text summarization in 2004. They compute sentence importance using a graph representation of sentences to extract salient textual units from documents as summarization. In 2020, Wang et al. propose to extract salient textual units from documents as summarization using a heterogeneous graph consisting of semantic nodes at several granularity levels of documents (Wang et al. 2020). In 2022, Wang et al. incorporate knowledge graphs into document encoding and decoding, generating the summary from a knowledge graph template to achieve state-of-the-art performance (Wang et al. 2022). However, to the best of our knowledge, no existing work integrates LLMs into graph-based methods to leverage their natural language understanding capabilities for improved graph construction and summary generation.
@@
In the first iteration, we apply the knowledge minigraph prompt to the first chunk of reference papers to derive the initial knowledge minigraph . For subsequent iterations, we transform the intermediate knowledge minigraph into a text representation . This text representation is then input along with the -th chunk of reference papers , to derive knowledge minigraph . After iterations, we obtain the final knowledge minigraph .
@@
Proposed | 36.41 | 8.78 | 34.16 | 6.22 | 32.31 | 5.36 |
Summarization router We design a router to evaluate different experts’ summaries and automatically select the most desirable summary as the final output. Without requiring additional side information, this paper proposes a self-evaluation strategy. In detail, we observe that there are agreements between the experts’ viewpoints and their generated summaries. We propose to quantify the degree of agreement for each summary using the ROUGE-1 score (Lin 2004), which measures the overlap between a generated summary and other summaries. We then select the summary with the highest degree of agreement, which indicates that its understanding has the highest likelihood of being supported by other experts, or in other words, is relatively more acceptable. Mathematically, the final summary is determined by:
@@
This section presents ablation studies to investigate the performance gains brought by the designed modules in CKMAs. We first ablate the knowledge minigraph construction agent (KMCA) and the multiple path summarization agent (MPSA), respectively. When ablating KMCA, we no longer construct the knowledge minigraph and do not include it as part of the MPSA instruction. When removing MPSA, we use the instruction in Table 1 to generate a single summary as the final summary directly. We report the performance of the ablated version in the first line of Table 3.